OpenStudy (anonymous):

solve for x: (9/7)log base 3 of x = -1

OpenStudy (zzr0ck3r):

\(\frac{9}{7}\log_3(x)=-1\implies\log_3(x)=-\frac{7}{9}\implies 3^{-\frac{7}{9}}=x\)

OpenStudy (zzr0ck3r):

\(\log_a(b)=c\iff a^c=b\). So first we just get the \(\log\) by itself, and then use that property. Does this make sense?